OpenStudy (anonymous):

which of the following is a polynominal with roots 2, -3i,and3i?

OpenStudy (anonymous):

the one that looks like \[(x-2)(x+3i)(x-3i)\] after you multiply out

OpenStudy (anonymous):

not as hard as it looks since \((x+3i)(x-3i)=x^2+9\) so really you have to multiply \[(x-2)(x^2+9)\]

OpenStudy (anonymous):

what i don't understand is the 3i.

OpenStudy (anonymous):

it is another name for \(3\sqrt{-1}\)

OpenStudy (anonymous):

oh. okay. i got it :) Thank you.

OpenStudy (anonymous):

\(i=\sqrt{-1}\) by definition so if you solve \[x^2+9=0\] you get \(x^2=-9\) or \(x=\pm3i\)

OpenStudy (anonymous):

yw